OpenStudy (icalibear):

How to solve systems with elimination? 5x - 2y = 28 -20x + 8y = -100 I know that if there are two of the same variable then they cancel and then it is easy to solve when you add the equations together, but what if neither equation has corresponding variables???

OpenStudy (anonymous):

you have to make them become corresponding 5x-2y= 28 (multiply by 8 ) = 80x-16y =224

OpenStudy (icalibear):

How are those corresponding :s

OpenStudy (anonymous):

and you have to make this for the second one -20x +8y=-100 (muliply by 2) = -40x +16y = -200 and solve by elimination now

OpenStudy (anonymous):

eliminate 16y

OpenStudy (icalibear):

why did you multoply by 8 and 2 sorry

OpenStudy (usukidoll):

If I was given that problem I would eliminate x... seems easier if I distribute the 4 all over the first equation

OpenStudy (usukidoll):

and for that reason. I have to distribute the 4 all over the first equation because I want to eliminate a variable and solve for the other. In this case, 4(5x-2y=28) 20x-8y=112 (this is my new first equation) so I add my new first equation to the second equation so I can get rid of the x

OpenStudy (anonymous):

to make them corresponding...

OpenStudy (anonymous):

@Michele_Laino please help

OpenStudy (icalibear):

i got 3/5...

OpenStudy (usukidoll):

oh geez we are forced to take the y's out and then solve for x because by distributing the 4 all over the first equation, I ended up destroying x and y., and we don't want that.

OpenStudy (usukidoll):

\[8(5x-2y=28)\] \[2(-20x+8y=-100)\] \[40x-16y=224\] \[-40x+16y=-200\] HUH? Am I reading this wrong? even I got the same problem.

OpenStudy (anonymous):

i was right @icalibear

OpenStudy (icalibear):



OpenStudy (icalibear):

i meant to type 40x

OpenStudy (icalibear):

sorry! typo

OpenStudy (icalibear):

i will fix this

OpenStudy (usukidoll):

we are supposed to distribute 8 and 2 THROUGHOUT THE equation. We can't have it one sided

OpenStudy (icalibear):